# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v2

@@ VI-G Time consumption
Table IV also lists the time consumption of ADOSD4 in terms of MBP4 iterations. Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations. This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions.
VI-G2 HRSR and OSD-0 in ADOSD4
We have shown that ADOSD4 is efficient, thanks to Algorithm 3 and Corollary 10. These methods reduce the effective length using reliable subset reduction and allow us to perform only OSD4-0 by identifying degenerate error cosets. Now, we examine the effectiveness of these techniques on the surface code of distance 11, with similar results observed for distances 13 and 15.
Table V enumerates the percentage of ADOSD4 trials that only perform OSD4-0 using Corollary 10. At $\epsilon=4.1\%$, 93.61% of ADOSD4 trials already perform only OSD4-0, increasing to 99.84% at $\epsilon=1.7\%$. Thus, at lower error rates, almost all trials execute OSD4-0.
Moreover, the HRSR algorithm reduces the dimensions of the check matrix in both rows and columns. Table V also provides the percentages of trials where the number of rows and columns in the check matrix is reduced to 20% or 30% of the original (denoted as $20\%$ and $30\%$ effective dimensions in the table). These percentages increase quickly as the error rate decreases. With fewer than 30% effective dimensions, the time consumption for Gaussian elimination reduces to roughly $0.3^{3}=2.7\%$, consistent with the previous results.
TABLE V: Percentage of OSD-0 in ADOSD4 on the surface code of distance $11$
depolarizing rate $\epsilon$ 1.7% 2.7% 4.1%
LER $10^{-6}$ $10^{-5}$ $10^{-4}$
OSD-0 99.84% 98.85% 93.61%
20% effective 44.13% 26.30% 5.13%
dimensions
30% effective 94.04% 84.46% 57.91%
dimensions
VII Conclusion
We proposed quaternary OSD with integrated hard- and soft-decision reliabilities to retain the $X/Z$ correlations. Our BP-OSD schemes have demonstrated highly effective decoding capabilities for various sparse quantum codes in the code capacity noise model, including both CSS and non-CSS codes. Additionally, when the logical error rate is around $10^{-6}$, ADOSD4 can further enhance the decoding process by offering accelerated decoding while maintaining comparable performance. As a result, ADOSD4 is applicable to small or medium-sized codes with $n\approx 1000$, significantly enhancing the performance of BP in the code capacity noise model. In our simulations, we restricted the number of error candidates to the size of order-2 OSD; better performance of ADOSD4 can be achieved by allowing for greater options.
We note that while all the techniques developed in this paper are presented in the context of quaternary error variables, they can naturally be adapted to binary OSD for handling CSS codes. This includes hard-decision reliability, highly reliable subset reduction, approximate degenerate decoding, and degenerate candidate pruning conditions, by restricting the approach to either $X$- or $Z$-stabilizers. Furthermore, these techniques can be integrated with other methods available in the literature.
In our investigations, we primarily focused on MBP4+OSD4-2 and ADOSD4. However, MBP4+OSD4-0 has also exhibited excellent performance for most codes, aligning with the findings in [43]. In particular, we found that ADOSD4 mostly performs only MBP4+OSD4-0 at low error rates. On the other hand, at low error rates (e.g., below $10^{-3}$), there are small non-isolated errors, so the effective length for OSD is relatively small, allowing us to perform higher-order OSD. We have observed some instances of surface codes with order $w$ up to 13 in (16).
An interesting observation is that we can identify bit-flips in error candidates corresponding to multiplying a stabilizer operator, as stated in Theorem 8 and Corollary 10. This information can be used to improve the algorithm by flipping effective bits in approximate degenerate decoding. Additionally, another potential improvement is extending the candidate pruning condition by Fossorier et al[47] to quantum codes.
We chose the binary representation of quantum codes and syndrome correction for developing the OSD algorithm. A quantum code can be interpreted as a quaternary additive code with binary error syndromes [6]. Thus, one might consider formulating a system consisting of $n$ quaternary variables and $n-k$ additive constraints and employ a Gaussian elimination process for quaternary additive groups. However, for practical implementation, we chose to binarize this process.
There are various reliability metrics that consider the (weighted) accumulation of the log-likelihood ratio (LLR) signs for each bit during BP iterations [72, 73, 74, 75]. Our hard-decision reliability can be viewed as an extension of this idea, focusing on the length of the last run in the hard-decision vectors. The reliability order in Definition 4 can be adapted to employ alternative metrics, such as Entropy or Max [43]. From our experiments, the integrated hard- and soft-decision reliabilities yielded the best results.
Finally, there is significant progress in circuit-level error decoding [76, 77, 78, 56, 79]. Various decoders have been shown to work for circuit-level noise models, including BP-OSD-CS, GDG, ambiguity clustering[57], LSD [58], and ordered Tanner forest [80]. Our BP-OSD scheme can be extended to handle both data and syndrome errors in the phenomenological noise model [81, 82], using techniques proposed in [83, 84, 85], as demonstrated in our recent work [86].

### v3

@@ VII-B1 Bivariate bicycle codes
Table III also lists the time consumption of ADOSD4 in terms of MBP4 iterations. Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations. This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions.
VII-C Circuit-Level Surface Code Decoding
We simulate circuit-level decoding of the $d^{2},1,d]]$ rotated surface codes using DEM matrices generated by STIM [71]. For a distance-$d$ code, we simulate $d$ rounds of syndrome extraction.
All faulty locations are assumed to have the same error rate $\epsilon$. We use stim.Circuit with experiment surface_code:rotated_memory_z to generate the syndrome-extraction circuits, and convert the resulting DEM into the binary matrices ${\mathbf{H}}_{\mathrm{DEM}}$, ${\mathbf{L}}_{\mathrm{DEM}}$, and the probability vector ${\mathbf{p}}_{\mathrm{DEM}}(\epsilon)$, where ${\mathbf{p}}_{\mathrm{DEM}}(\epsilon)$ depends on the physical error rate $\epsilon$.
In the default setting, the first round of detector variables contains detectors for only the $\frac{d^{2}-1}{2}$ $Z$ stabilizers. The subsequent d1d-1 rounds contain detector variables for all $d^{2}-1$ $X$ and $Z$ stabilizers. In addition, there is a final round of detector variables for the $Z$ stabilizers, which can be regarded as noiseless syndrome measurements. As a result, the default STIM circuit for a $d^{2},1,d]]$ rotated surface code yields a total of $d(d^{2}-1)$ parity checks.
Note that decoders which treat $X$ and $Z$ errors separately, such as the standard MWPM decoder, do not make use of both types of detector variables. For comparison, we therefore also generate decoding problems in which the $X$-type detector variables are removed. The parameters of the resulting DEM decoding problems are summarized in Table IV.
Since the default DEM contains more detector variables and hence more syndrome bits, it produces a larger number of error classes, i.e., a larger $N$, because some $Y$ errors can be distinguished from $Z$ errors by the presence of $X$-type detector variables.
Distance 3 5 7 9
NN 219 1677 5471 12705
# Checks 24 120 336 720
(a) DEM decoding parameters with both $X$ and $Z$ detector variables.
Distance 3 5 7 9
NN 55 301 883 1945
# Checks 16 72 192 400
(b) DEM decoding parameters with only $Z$ detector variables.
TABLE IV: Parameters of the DEM decoding problems for rotated surface codes. Subtable (a) shows the default setting with both $X$ and $Z$ detector variables, and subtable (b) shows the modified setting with only $Z$ detector variables.
In our simulations, an error vector $\mathbf{e}$} is sampled according to ${\mathbf{p}}_{\mathrm{DEM}}(\epsilon)$ and its syndrome ${\mathbf{H}}_{\mathrm{DEM}}{\mathbf{e}}^{\top}$ is computed. Given ${\mathbf{H}}_{\mathrm{DEM}}$, ${\mathbf{L}}_{\mathrm{DEM}}$, the measured syndrome ${\mathbf{H}}_{\mathrm{DEM}}{\mathbf{e}}^{\top}$, and ${\mathbf{p}}_{\mathrm{DEM}}(\epsilon)$, we apply MBP4+ADOSD4 to this binary decoding problem to obtain an estimate $\hat{{\mathbf{e}}}$. The logical $X$ error rate is then estimated using (21) and (22).
We collected 1000 logical error events for each data point in the plots, except for the cases with $d=9$ and $N=12705$ at $\epsilon=0.002$ and $\epsilon=0.001$.
VII-C1 Decoding performance
For computational efficiency, the maximum number of BP iterations is set to $T=10$. In particular, we use only soft reliability with threshold $\theta=0.99$, without employing hard-decision history. By lowering the reliability threshold, RSR produces a smaller reduced decoding problem.
The step size is fixed to $\alpha=1.5$, following the choice used in the code-capacity noise model without further optimization.
Figure 7: MBP4+ADOSD4 and MWPM decoding performance for various $d^{2},1,d]]$ rotated surface codes with $\alpha=1.5$, $T=10$, and $\theta=0.99$ under the circuit-level noise model. Error bars represent 95% confidence intervals.
The simulation results are shown in Fig. 7, where the MWPM decoder is implemented using PyMatching2 [89] for comparison. We observe that MWPM performs better than MBP4+ADOSD4 for DEM instances without $X$-type detector variables. However, when $X$ detector variables are included, MBP4+ADOSD4 outperforms MWPM by exploiting the additional syndrome information. Moreover, for higher physical error rates ($p>0.003$), the $d=9$ MBP4+ADOSD4 curve outperforms the $d=11$ MWPM curve. The error threshold is also improved to approximately $0.76\%$, compared with the $0.7\%$ reported for PyMatching2.
For reference, as shown in Ref. [90], RelayBP performs worse than MWPM for rotated surface codes. We also remark that MWPM can be adapted to exploit $X/Z$ correlations [91, 92, 93].
Furthermore, we include a direct comparison with BP+LSD-0 from [55] in Fig. 8, where our decoder consistently achieves a lower logical error rate. Although it is unclear whether [55] uses the default DEM generated by stim, both of our performance curves (with and without $X$ detector variables) outperform BP+LSD-0.
Figure 8: MBP4+ADOSD4 and BP–LSD-0 decoding performance for $d^{2},1,d]]$ rotated surface codes with $\alpha=1.5$, $T=10$, and $\theta=0.99$ under the circuit-level noise model. The BP+LSD-0 curves are from [55] and the algorithm uses a maximum of 30 BP iterations.
VII-C2 Decoding statistics
We summarize the decoding statistics for circuit-level surface-code decoding of distance $9$ without $X$-type detector variables in Table V. The row “Samples” reports the total number of Monte Carlo shots, while “OSD calls” counts the number of samples for which BP does not converge to a valid solution and OSD is therefore invoked.
The row “Average length after RSR” shows that the effective code length after reduction is below $5\%$ of the original for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$, demonstrating the strong reduction capability of RSR.
From the statistics, we observe no stage-1 failures and only a small number of stage-2 failures. Even if all stage-2 failures are conservatively counted as logical errors, the overall logical error rate is only marginally affected.
The row “Avg iterations w/o OSD” corresponds to samples that are successfully decoded by BP alone without invoking OSD. Typically, only about 3366 BP iterations are required. This justifies setting the maximum number of BP iterations to $T=10$, which is sufficient for RSR to accurately identify reliable and unreliable bits.
“Time per BP iteration” denotes the average time of a single BP iteration, averaged over all runs. “Time per OSD” is the average time of one OSD call, including both ADOSD and the fallback regular OSD when RSR fails. We observe that the time per BP iteration is essentially independent of the depolarizing rate $\epsilon$. In contrast, the OSD time decreases significantly as $\epsilon$ becomes smaller. At $\epsilon=0.01$, one OSD call is equivalent to roughly 250 BP iterations, while at $\epsilon=0.001$ it drops to about 18 BP iterations.
For reference, in the reported simulations of Relay-BP-5, $R=601$ and $T_{r}=60$, resulting in more than $3.6\times 10^{4}$ BP iterations [90].
Finally, we remark that the current OSD runtimes do not yet fully reflect the potential speedup enabled by RSR. Our implementation embeds the binary DEM decoding problem into a quaternary formalism, which is not optimized for this task and can be further accelerated by a dedicated binary implementation.
TABLE V: Statistics of circuit-level surface decoding of distance $9$ without $X$ detector variables. $N=1945$.
depolarizing rate $\epsilon$ 0.01 0.007 0.005 0.003 0.002 0.001
Samples 100,000 100,000 115990 1311328 9943799 10,000,000
OSD calls 99966 96895 91954 483438 1632156 389602
Average length after RSR 229.76 113.37 62.64 32.69 22.67 14.99
relative code length 0.12 0.06 0.03 0.02 0.012 0.008
Stage-1 Failures 0 0 0 0 0 0
Stage-2 Failures 15 43 109 1142 3927 916
OSD-0 0 0 0 2 107 219
Avg iterations w/o OSD 5.824 5.691 5.144 4.525 4.059 3.353
Time per BP iteration (ms) 0.1565 0.1498 0.1289 0.1248 0.1219 0.1185
Time per OSD (ms) 37.65 27.04 20.71 19.90 8.632 1.987
VIII Related Work
Our decoder essentially follows a BP+RSR+OSD pipeline: 1) run BP; 2) if BP does not converge to a valid codeword, use the reliability statistics generated by BP to perform RSR and obtain a reduced linear system; 3) apply OSD to solve the reduced system. In this way, the postprocessing stage is significantly accelerated.
Related work such as Ambiguity Clustering (AC) [54] and Localized Statistics Decoding (LSD) [55] also aim to speed up the postprocessing stage. AC employs a divide-and-conquer strategy by decomposing the global decoding problem into multiple localized clusters. Each cluster is solved independently to determine its local logical effect, and the local solutions are then linearly combined to reconstruct a global logical correction.
LSD follows a cluster-growth paradigm inspired by the Union-Find (UF) decoder [94] to localize syndromes. It further uses linear-algebraic verification to govern cluster merging and syndrome satisfaction, allowing LSD to be applied to codes without a geometric structure.
Since AC and LSD can be viewed as high-efficiency alternatives to OSD for postprocessing, it is natural to integrate them into a unified pipeline such as BP+RSR+AC or BP+RSR+LSD. Note that RSR requires performing rank-revealing linear-algebraic operations on the reduced system. Importantly, essentially the same class of linear-algebraic operations (e.g., pivoting, rank tests, and column selections) is also required by OSD, AC, and LSD in order to initialize clusters, perform algebraic verification, or generate candidate solutions. Therefore, applying AC or LSD after RSR does not introduce a fundamentally new computational stage.
We further note that AC applies degenerate decoding when solving the local subproblems. Because these subproblems are much smaller, it becomes feasible to perform fully degenerate decoding by optimizing over logical cosets. This observation motivates the use of $\delta$-ADD with $\delta>0$ as a hard-decision rule in OSD after RSR has reduced the problem size sufficiently.
To summarize, degeneracy is exploited in our algorithm in two distinct ways. (1) Since OSD-$w$ is a list decoder, we first use a degeneracy condition (Theorem 8 and Corollary 9) to determine whether multiple logical cosets are possible, and hence whether generating a candidate list is necessary at all. (2) Once a reduced problem has been obtained by RSR and a candidate list is generated, we apply the approximate degenerate decoding rule $\delta$-ADD to select the most likely logical coset from the OSD list.
IX Conclusion
We proposed an RSR procedure that reduces the decoding problem size using BP-generated statistics with both hard- and soft-decision reliabilities. RSR serves as a general-purpose preprocessor that enables post-processing decoders, including OSD, LSD, and AC, to operate efficiently on the reduced problems.
We further introduced a degeneracy-aware pruning condition to identify when higher-order OSD is beneficial. Combined with MBP, this yields the MBP–RSR–ADOSD family of decoders, which achieves strong and consistent performance across a broad range of sparse quantum codes under the code-capacity noise model, including both CSS and non-CSS codes, as well as under circuit-level DEM noise for rotated surface codes.
A key practical result is that MBP–RSR–ADOSD4 scales to decoding problems with more than $10^{4}$ error variables. At low physical error rates (e.g., below $10^{-3}$), errors are typically sparse and isolated, so RSR produces a small effective problem size, making higher-order OSD both computationally feasible and highly effective in this regime.
More general circuit-level decoding problems with strongly correlated syndrome structures [70] are left for future work.
The source files of our simulations can be found in [95].